OpenStudy (gabylovesyou):

@jim_thompson5910 What Is the equation of a line (in standard form) that passes through (-3,0) and (-1,-4) I got 2x + y = 6 correct ?

jimthompson5910 (jim_thompson5910):

incorrect 2x + y = 6 2(-3) + 0 = 6 ... replace (x,y) with (-3,0) -6 + 0 =6 -6 = 6 ... this is false, so (-3,0) is NOT on the line 2x+y = 6

OpenStudy (gabylovesyou):

ok idk where i went wrong.. first is slope which i got -2 correct ?

OpenStudy (gabylovesyou):

then, point-slope form and i got y - 0 = -2(x - -3)

OpenStudy (gabylovesyou):

then slope intercept form and i got y = -2x + 6

jimthompson5910 (jim_thompson5910):

well y - 0 = -2(x - (-3)) turns into y = -2(x + 3) and that turns into y = -2x - 6

OpenStudy (gabylovesyou):

wait why does it turn into that

jimthompson5910 (jim_thompson5910):

because the two negatives make a positive

jimthompson5910 (jim_thompson5910):

which is why y - 0 = -2(x - (-3)) turns into y = -2(x + 3)

jimthompson5910 (jim_thompson5910):

then you distribute it to y = -2x - 6

OpenStudy (gabylovesyou):

ok so then... what do i do ?

jimthompson5910 (jim_thompson5910):

you add 2x to both sides y = -2x - 6 y+2x = -2x - 6+2x 2x+y = -6

jimthompson5910 (jim_thompson5910):

I recommend you check both points in 2x+y = -6

OpenStudy (gabylovesyou):

and thats the standard form ?

jimthompson5910 (jim_thompson5910):

yes it's in Ax + By = C form

OpenStudy (gabylovesyou):

Thank you

jimthompson5910 (jim_thompson5910):

np